OpenStudy (anonymous):

Without using a calculator determine the exact value of sin(x - y) given that sin x = -1/3) , tan y = (1/3) , x is a quadrant II angle and y is a quadrant III angle.

OpenStudy (anonymous):

sin(x) shouldn't be negative if x is in the second quadrant. Is there a typo?

OpenStudy (anonymous):

Im sorry the problem is Without using a calculator determine the exact value of cos(x - y) given that sin x = (1/3) , tan y = (1/3) , x is a quadrant II angle and y is a quadrant III angle.

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

use sum/diff of angle expansion

OpenStudy (anonymous):

= cos(x)cos(y)+sin(x)sin(y)

OpenStudy (anonymous):

if sin(x) = (1/3) , draw up a general right angle triangle , mark an angle x ( obviously not the right angle ) , and label the opposite 1 and hypotenuse 3

OpenStudy (anonymous):

then you can find the adjacent side by pythagorus , which works out as sqrt(8) therefore cos(x) = - sqrt(8) / 3 ( because x is negative in the 2nd quadrant

OpenStudy (anonymous):

now you just need to find siny and cosy

OpenStudy (anonymous):

Possible answers are \[A. 0 B. (12\sqrt{5} - \sqrt{10}) /30 C. (-12\sqrt{5} - \sqrt{10}) /30 D. (-12\sqrt{5} + \sqrt{10}) /30\] \[E. (12\sqrt{5} + \sqrt{10}) /30\]

OpenStudy (anonymous):

with tan(y)=(1/3) , draw up a general right angle triangle , mark opposite as 1 and adjacent as 3

OpenStudy (anonymous):

therefore the hypotenuse is sqrt(10) so sin(y) = - 1 / sqrt(10) and cosy =- 3/sqrt(10)

OpenStudy (anonymous):

then sub in

OpenStudy (anonymous):

So my equation should look like (-sqrt(8)/3 * -3/sqrt(10) ) + (1/3 * -1/sqrt(10)) ???